OpenStudy (anonymous):

show your work 7 + 4x - 9 = -6

OpenStudy (merchandize):

4x=-6+9-7 4x=-4 x=-1

OpenStudy (anonymous):

7+4x-9=-6 7-9+4x=-6 -2+4x=-6 4x=-4 x=-1